OpenStudy (anonymous):

Find the limit: lim((x+2)/(x-3))^x as x approaches to positive infinity. Find the limit: lim((x+2)/(x-3))^x as x approaches to positive infinity. @Mathematics

OpenStudy (anonymous):

e^5

OpenStudy (turingtest):

\[\lim_{x \rightarrow \infty}({x+2\over x-3})^x=\infty\]is all I see. How did you get your answer moneybird?

OpenStudy (anonymous):

moneybird can you write down the process ?

OpenStudy (turingtest):

yes please

OpenStudy (zarkon):

moneybird is correct

OpenStudy (anonymous):

can you write down the process Zarkon? because i want to know how.

OpenStudy (zarkon):

can you use \[\lim_{x\to\infty}\left(1+\frac{y}{x}\right)^x=e^y\] ?

OpenStudy (anonymous):

why do i use that form?

OpenStudy (zarkon):

if you know the above then you can use it to find the answer to your problem... otherwise you can use \[\left({x+2\over x-3}\right)^x=\exp\left(x\ln\left({x+2\over x-3}\right)\right)\] \[=\exp\left(\frac{\ln\left({x+2\over x-3}\right)}{1/x}\right)\] now use L'Hospital's rule

OpenStudy (anonymous):

I am stuck with using the rule. Could you tell me how to solve?

OpenStudy (zarkon):

write \[\ln\left({x+2\over x-3}\right)=\ln(x+2)-\ln(x-3)\] in the above function then use L'Hospital's rule.

OpenStudy (anonymous):

oh okay.